The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall common stock was $ 64 on December 31, 20Y2.

Marshall Inc.
Comparative Retained Earnings Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Retained earnings, January 1 $ 4,159,100 $ 3,518,300
Net income 921,600 720,600
Total $5,080,700 $ 4,238,900
Dividends:
On preferred stock $ 11,900 $ 11,900
On common stock 67,900 67,900
Total dividends $ 79,800 $ 79,800
Retained earnings, December 31 $ 5,000,900 $ 4,159,100


Marshall Inc.
Comparative Income Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Sales $ 5,450,910 $ 5,022,190
Cost of goods sold 1,949,100 1,793,170
Gross profit $ 3,501,810 $ 3,229,020
Selling expenses $ 1,185,450 $ 1,442,200
Administrative expenses 1,009,830 847,010
Total operating expenses $2,195,280 $2,289,210
Income from operations $ 1,306,530 $ 939,810
Other revenue 68,770 59,990
$ 1,375,300 $ 999,800
Other expense (interest) 328,000 180,800
Income before income tax $ 1,047,300 $ 819,000
Income tax expense 125,700 98,400
Net income $ 921,600 $ 720,600


Marshall Inc.
Comparative Balance Sheet
December 31, 20Y2 and 20Y1
   20Y2    20Y1
Assets
Current assets
Cash $ 921,420 $ 974,710
Marketable securities 1,394,590 1,615,230
Accounts receivable (net) 985,500 927,100
Inventories 730,000 569,400
Prepaid expenses 174,322 194,940
Total current assets $ 4,205,832 $ 4,281,380
Long-term investments 3,151,788 1,733,000
Property, plant, and equipment (net) 4,920,000 4,428,000
Total assets $ 12,277,620 $ 10,442,380
Liabilities
Current liabilities $ 1,356,720 $ 2,203,280
Long-term liabilities:
Mortgage note payable, 8% $ 1,840,000 $ 0
Bonds payable, 8% 2,260,000 2,260,000
Total long-term liabilities $ 4,100,000 $ 2,260,000
Total liabilities $ 5,456,720 $ 4,463,280
Stockholders' Equity
Preferred $0.70 stock, $50 par $ 850,000 $ 850,000
Common stock, $10 par 970,000 970,000
Retained earnings 5,000,900 4,159,100
Total stockholders' equity $ 6,820,900 $ 5,979,100
Total liabilities and stockholders' equity $ 12,277,620 $ 10,442,380

Required:

Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year.

1. Working capital $
2. Current ratio
3. Quick ratio
4. Accounts receivable turnover
5. Number of days' sales in receivables days
6. Inventory turnover
7. Number of days' sales in inventory days
8. Ratio of fixed assets to long-term liabilities
9. Ratio of liabilities to stockholders' equity
10. Times interest earned
11. Asset turnover
12. Return on total assets %
13. Return on stockholders’ equity %
14. Return on common stockholders’ equity %
15. Earnings per share on common stock $
16. Price-earnings ratio
17. Dividends per share of common stock $
18. Dividend yield %
Answer #1
20Y2
1 Working Capital Current Asset - Current Liabilities $2849112 =4205832-1356720
2 Current ratio Current Assets / Current Liabilities 3.1 =4205832/1356720
3 Quick Ratio (Current Assets - Inventory - Prepaid Expenses )/Current Liabilities 2.4 =(4205832-730000-174332)/1356720
4 Accounts Receivables turnover Net Credit Sales / Average Accounts receivable 1.4 =5450910/(985500+927100)/2
5 Number of days' sales in receivables 365 days / Receivables turnover ratio 260.7 Days =365/1.4
6 Inventory Turnover Cost of goods sold / average inventory 0.2 =1949100/(730000+5694000)/2
7 Number of days' sales in inventory 365 days / inventory turnover ratio 182.5 Days =365/2
8 Ratio of fixed assets to long-term liabilities Total Fixed Assets / Long term Liabilities 2.0 =(4920000+3151788)/4100000
9 Ratio of liabilities to stockholders' equity Total liabilities /Shareholder's equity 0.8 =5456720/6820900
10 Times interest earned Earnings before interest and taxes / total interest 4.2 Times =1375300/328000
11 Asset turnover Net sales / average total assets 0.1 =5450910/(12277620+10442380)/2
12 Return on total assets Net income / total assets 7.5% =921600/12277620
13 Return on Stockholders equity Net income / Stockholders equity 13.5% =921600/6820900
14 Return on common stockholders’ equity Net income / Common shareholder's equity 15.4% =921600/(6820900-850000)
15 Earnings per share (Net Income - Preferred Dividend )/ Total no of common shares $9.38 =(921600-11900)/97000
16 Price-earnings ratio Current share price / Earnings per share                          6.8 =64/9.38
17 Dividend per share on common stock Dividend to common stock holders / Total number of common stock $0.70 =67900/97000
18 Dividend yield Dividend Per share/ Current Share Price 1.1% =0.7/64